ORTHO-PARA MIXING INTERACTION IN THE VINYL RADICAL DETECTED BY MILLIMETER-WAVE SPECTROSCOPY AND PREDICTION OF FAST ORTHO-PARA CONVERSION RATE

\underline{KEIICHI TANAKA}, MASATO HAYASHI, MITSUHIKO OHTSUKI, KENSUKE HARADA, and TAKEHIKO TANAKA, Department of Chemistry, Faculty of Science, Kyushu University, Hakozaki, Higashiku, Fukuoka 812-8581, Japan.

Ortho-para mixing interaction due to the coupling of nuclear and electron spins was detected for the first time by millimeterwave spectroscopy of deuterated vinyl radicals, H_{2}CCD and D_{2}CCD, of which the ground states are split by the tunneling motion of the α deuteron into two components 0^{+} and 0^{−}, whose separations have been determined to be ΔE_{0} = 1186.644(16) and 771.978(18) MHz, respectively. The observed tunneling-rotation spectra are significantly perturbed by the othro-para mixing interaction expressed by ⟨0^{±}|H′|0^{∓}⟩ = (δa_{F}^{(β)}/2)S · (I_{β1} − I_{β2}), where I_{β1} and I_{β2} are spins of the two hydrogen nuclei in the β position and S is the electron spin, which connects rotational levels in the 0^{+} and 0^{−} states, one being an ortho level and the other a para level. The δa_{F}^{(β)} constants for H_{2}CCD and D_{2}CCD have been determined to be 68.06(53) and 10.63(94) MHz, respectively, consistent each other within the isotopic mass relation. The othro and para states are mixed by about 0.097% and 0.0123% due to this interaction. ^{a} The δa_{F} constant for H_{2}CCH should be similar to that for H_{2}CCD because of the same probability density of the unpaired electron at the β protons, but could not be determined independently in our previous study. It is because the mixing of para- and ortho-levels of about 0.00044% is much smaller than that for H_{2}CCD due to the large tunneling splitting of ΔE_{0} = 16271.8429(59) MHz. ^{b} The rate constant of para to ortho (I_{β} = 0 → 1) conversion is predicted as 1.2 × 10^{5} s^{−1} torr^{−1} for H_{2}CCD, suggesting extremely rapid mutual conversion between ortho and para nuclear spin isomers of H_{2}CCD, which is more than 10^{6} times faster compared with that in closed shell molecules such as H_{2}CO and H_{2}CCH_{2}.

aK. Tanaka, M. Hayashi, M. Ohtsuki, K. Harada, T. Tanaka, J. Chem. Phys. \underline{131,} 111101 (2009).

bK. Tanaka, M. Toshimitsu, K. Harada, T. Tanaka, J. Chem. Phys. \underline{120,} 3604 (2004).